# Three Score and 15 Years (1948-2023) of Rao’s Score Test: A Brief History

### v2

@@ Abstract
2.1.From Rao and Poti, (1946) to Rao, (1948): Test for the Multiparameter Case
@@
Example 1. Let us start with a simple example where $y_{i}\sim IIDN(\theta,1),\ i=1,2,\ldots,n$, and we test $H_{0}:\theta=\theta_{0}=0$ against $\theta>0$. Here the log-likelihood and score functions are respectively
@@
Example 2. [Ferguson, (1967, p.235)] Consider testing for the median of a Cauchy distribution with density
@@
Example 3. [Pearson, (1900) Goodness-of-fit test]. Consider a multinomial distribution with $p$ classes and let the probability of an observation belonging to the $j$-th class be $\theta_{j}(\geq 0)$, $j=1,2,\ldots,p$, so that $\sum^{p}_{i=1}\theta_{j}=1$. Denote the observed frequency of the $j$-th class by $n_{j}$ with $\sum^{p}_{j=1}n_{j}=n$. We are interested in testing $\theta_{j}=\theta_{j0}$, $j=1,2,\ldots,p$, where $\theta_{j0}$ are known constants. Pearson, (1900) suggested the statistic
@@
Example 5.1 Suppose we take $f(y;\theta)\equiv N(\mu,\sigma^{2})$, and let the DGP $g(y)$ satisfy $E_{g}[y]=\mu$, $E_{g}[y-\mu]^{2}=\sigma^{2}$, $E_{g}[y-\mu]^{3}=\mu_{3}$ and $E_{g}[y-\mu]^{4}=\mu_{4}$. Then it is easy to show that
@@
Example 5.2 In (37) the parameters $c_{1}$ and $c_{2}$ of the Pearson family of distributions can be treated, respectively, as the “skewness” and “kurtosis” parameters. Suppose we test the symmetry ignoring the (excess) kurtosis. Then we can start with the system (37) with $c_{2}=0$, that is,
@@
Example 5.3 Consider the test for homoskedasticity under the regression framework of (36), where we now explicitly specify the heteroskedastic structure as $Var(\epsilon_{i})=\sigma^{2}_{i}=\sigma^{2}+\delta^{\prime}z_{i}$, where $\delta$ is a $r\times 1$ vector and $z_{i}$’s are fixed exogenous variables, $i=1,2,\ldots,n$. Assuming normality of $\epsilon_{i}$ the RS statistic for testing homoskedasticity hypothesis $H_{0}:\delta=0$ is given by [see Breusch and Pagan, (1979)]
@@
Example 5.4 To illustrate the usefulness of the robust score statistic $RS^{*}_{\psi}(P)$, we now consider the tests developed in Anselin et al., (1996) for the mixed regressive - spatial autoregressive (SAR) model with a SAR disturbance
@@
5.3.Robust RS Tests Under Both the Distributional and Parametric Misspecifications
@@
Example 5.5 Let us briefly go back to Example 5.4 and now introduce distributional misspecification along with the presence of parametric misspecification. This case has been rigorously considered by Fang et al., (2014) and they demostrated both analytically and through extensive simulations that $RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(P)$ as given, respectively in (74) and (77) are valid under non-normality. Therefore, $RS^{*}_{\psi}(DP)=RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(DP)=RS^{*}_{\phi}(P)$. This is a somewhat unsual situation. For this model as given in (72), information matrix equality does not hold, i.e., $J(\theta_{g})\neq K(\theta_{g})$ [see equations (51)-(53)]. However, still $J^{-1}KJ^{-1}=J^{-1}$. This is a serendipitous situation, since no additional adjustment is needed for the distributional misspecification. The intuition behind this serendipity is that the hypotheses $\psi=0$ and $\phi=0$ relate to the conditional mean (first moment) of $y$ in (72) (conditional on the neighborhood as captured by the $W$ matrix). However, in general, only tests for variance (second moment) and higher moments get affected by non-normality. A similar case appeared in Bera et al., (2020) where they considered testing for random effects and serial correlation within an error component model. Extensive simulation results are also given in Koley and Bera, (2022, 2024) demonstrating the robustness of the RS tests under non-normality in finite sample in spatial regression model set up.
@@ Acknowledgements
We are profoundly thankful to an anonymous referee for her/his careful reading and offering many pertinent comments which led to improvement of the paper. We are grateful to our research assistant (RA) Anirudh Adhikary. Without his assistance this paper wouldn’t have taken off. We are also grateful to our other RAs, Rong Yuwen, Tiancheng Guo, Scarlett He, Yice Zhang and Wenqi Zeng for their painstaking work on compiling the citation numbers and a very careful analysis. We wish we could report all the results they obtained. Thanks are also due to Dr. Malabika Koley for comments on an earlier draft, that improved the exposition of the paper. We however, retain the responsibility of any remaining errors. Part of this work was completed when the second author visited University of Illinois at Urbana-Champaign (UIUC).

### v3

@@ Abstract
2.1.From Rao and Poti, (1946) to Rao, (1948): Test for the multiparameter
case
@@
Example 1: Let us start with a simple example where $y_{i}\sim IIDN(\theta,1),\ i=1,2,\ldots,n$, and we test $H_{0}:\theta=\theta_{0}=0$ against $\theta>0$. Here the log-likelihood and score functions are respectively
@@
Example 2: [Ferguson, (1967, p.235)] Consider testing for the median of a Cauchy distribution with density
@@
Example 3: [Pearson, (1900) Goodness-of-fit test]. Consider a multinomial distribution with $p$ classes and let the probability of an observation belonging to the $j$-th class be $\theta_{j}(\geq 0)$, $j=1,2,\ldots,p$, so that $\sum^{p}_{i=1}\theta_{j}=1$. Denote the observed frequency of the $j$-th class by $n_{j}$ with $\sum^{p}_{j=1}n_{j}=n$. We are interested in testing $\theta_{j}=\theta_{j0}$, $j=1,2,\ldots,p$, where $\theta_{j0}$ are known constants. Pearson, (1900) suggested the statistic
@@
Example 4: Suppose we take $f(y;\theta)\equiv N(\mu,\sigma^{2})$, and let the DGP $g(y)$ satisfy $E_{g}[y]=\mu$, $E_{g}[y-\mu]^{2}=\sigma^{2}$, $E_{g}[y-\mu]^{3}=\mu_{3}$ and $E_{g}[y-\mu]^{4}=\mu_{4}$. Then it is easy to show that
@@
Example 5: In (37) the parameters $c_{1}$ and $c_{2}$ of the Pearson family of distributions can be treated, respectively, as the “skewness” and “kurtosis” parameters. Suppose we test the symmetry ignoring the (excess) kurtosis. Then we can start with the system (37) with $c_{2}=0$, that is,
@@
Example 6: Consider the test for homoskedasticity under the regression framework of (36), where we now explicitly specify the heteroskedastic structure as $Var(\epsilon_{i})=\sigma^{2}_{i}=\sigma^{2}+\delta^{\prime}z_{i}$, where $\delta$ is a $r\times 1$ vector and $z_{i}$’s are fixed exogenous variables, $i=1,2,\ldots,n$. Assuming normality of $\epsilon_{i}$ the RS statistic for testing homoskedasticity hypothesis $H_{0}:\delta=0$ is given by [see Breusch and Pagan, (1979)]
@@
Example 7: To illustrate the usefulness of the robust score statistic $RS^{*}_{\psi}(P)$, we now consider the tests developed in Anselin et al., (1996) for the mixed regressive - spatial autoregressive (SAR) model with a SAR disturbance
@@
5.3.Robust RS tests under both the distributional and parametric misspecifi-
cations
@@
Example 8: Let us briefly go back to Example 7 and now introduce distributional misspecification along with the presence of parametric misspecification. This case has been rigorously considered by Fang et al., (2014) and they demostrated both analytically and through extensive simulations that $RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(P)$ as given, respectively in (74) and (77) are valid under non-normality. Therefore, $RS^{*}_{\psi}(DP)=RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(DP)=RS^{*}_{\phi}(P)$. This is a somewhat unsual situation. For this model as given in (72), information matrix equality does not hold, i.e., $J(\theta_{g})\neq K(\theta_{g})$ [see equations (51)-(53)]. However, still $J^{-1}KJ^{-1}=J^{-1}$. This is a serendipitous situation, since no additional adjustment is needed for the distributional misspecification. The intuition behind this serendipity is that the hypotheses $\psi=0$ and $\phi=0$ relate to the conditional mean (first moment) of $y$ in (72) (conditional on the neighborhood as captured by the $W$ matrix). However, in general, only tests for variance (second moment) and higher moments get affected by non-normality. A similar case appeared in Bera et al., (2020) where they considered testing for random effects and serial correlation within an error component model. Extensive simulation results are also given in Koley and Bera, (2022, 2024) demonstrating the robustness of the RS tests under non-normality in finite sample in spatial regression model set up.
@@ Acknowledgements
We are profoundly thankful to an anonymous referee for her/his careful reading and offering many pertinent comments which led to improvement of the paper. An earlier version of the paper was presented at the Invited Memorial Session for Professor C. R. Rao, Joint Statistical Meetings (JSM), Portland, Oregon, August 3-8, 2024. We are thankful to the participants of that conference for their comments, especially to the organizer of the Session, Professor Ronald L. Wasserstein, Executive Director of the American Statistical Association (ASA), for giving us the opportunity to present our paper. This paper was also presented at the Department of Statistics, University of Illinois at Urbana-Champaign (UIUC). We would like to thank the attendees of the UIUC seminar for their constructive feedback that further helped in producing an improved version. We are grateful to our research assistant (RA) Anirudh Adhikary. Without his assistance this paper wouldn’t have taken off. We are also grateful to our other RAs, Rong Yuwen, Tiancheng Guo, Scarlett He, Yice Zhang and Wenqi Zeng for their diligent work on compiling the citation numbers and a very careful analysis. We thank Professor Osman Doğan for reading an earlier version of the paper with great care and thoroughness, and providing many pertinent comments with detail suggestions for improvements. We wish we could incorporate all his suggestions. Thanks are also due to Dr. Malabika Koley for comments that improved the exposition of the paper. At the publication stage, the Chair Editor, Professor V. K. Gupta paid painstaking attention to the finest details. Indeed, his suggestions, we believe, greatly improved the final presentation. We, however, retain the responsibility of any remaining errors. Part of this work was completed when the second author visited UIUC.